OpenStudy (jojo12g):

Which is a solution to the equation 5x – y = –3? A. (–2, 7) B. (2, –7) C. (1, 2) D. (–2, –7)

OpenStudy (jojo12g):

@Mehek14

Mehek (mehek14):

You can plug in each ordered pair

OpenStudy (jojo12g):

i think A

Mehek (mehek14):

Lets check 5 * -2 - 7 = -3 -10 - 7 = -3 In correct

OpenStudy (jojo12g):

so then D

Mehek (mehek14):

5 * - 2 = -10 -10 + 7 = -3 Yes

OpenStudy (jojo12g):

thank so it is D

Mehek (mehek14):

Yep